OpenStudy (friendzone):

2(x + 7) + 3x = 12 What property is used in the first step in solving this equation for x? A) Reflexive Property B) Transitive Property C) Commutative Property D) Distributive Property

OpenStudy (anonymous):

Distributive

OpenStudy (anonymous):

2(x+7)+3x = 12, now apply distibutive in the parenteses: 2x + 14 + 3x = 12 5x = 12-14 x = -2/5